OpenStudy (anonymous):

Having survived the meteor impact, thanks to some last-minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y2 + x2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’s path.

OpenStudy (anonymous):

original trajectory: y=6x+2

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (danjs):

hi

OpenStudy (anonymous):

hey,wats up? :)

OpenStudy (danjs):

So there is going to be a point where that line crosses the elipse

OpenStudy (anonymous):

yea i just dont know how to put my original trac. into the new problem? any ideas?

OpenStudy (danjs):

yeah...

OpenStudy (danjs):

put y = 6x+2 into the circle equation

OpenStudy (danjs):



OpenStudy (anonymous):

i am going to send to what i had got, give me a minute

OpenStudy (danjs):

ok, ill start typing out stuff

OpenStudy (anonymous):

y2+x2=40000 ,y=6x+2 6x2+2+x2=40000 7x2+2=40000 7x2=39998 I get stuck here.

OpenStudy (danjs):

\[(6x+2)^2 +x^2 = 40000\] \[[36x^2 + 24x + 4] + x^2 = 200^2\] \[37x^2 + 24x - 39996 = 0\]

OpenStudy (danjs):

remember you are squaring a quantity (a +b)^2 = a^2 +2ab + b^2

OpenStudy (danjs):

see that?

OpenStudy (anonymous):

ohh, okay so that's were i went wrong.

OpenStudy (danjs):

\[Ax^2 + Bx^2 + C = 0\] \[x = \frac{ -B \pm \sqrt{B^2 - 4*A*C} }{ 2*A }\]

OpenStudy (anonymous):

okay finna plug hem in an solve

OpenStudy (anonymous):

i mean them,lol

OpenStudy (danjs):

yep

OpenStudy (danjs):

ill start typing it up while you work at it

OpenStudy (danjs):

\[x = \frac{ -24 \pm \sqrt{24^2 - 4*37*(-39996)} }{ 2*37 }\] \[x = \frac{ -24 \pm \sqrt{5919984} }{ 74 }\] \[x = \frac{ -24 \pm 84\sqrt{839} }{ 74 } = \frac{ -24 }{ 74 } \pm \frac{ 84\sqrt{839} }{ 74 }\] \[x = \frac{ 12 }{ 37 } \pm \frac{ 42\sqrt{839} }{ 37 }\] \[x \approx 32.56 ~~or~~x \approx -33.20\]

OpenStudy (anonymous):

sorry, but i have to leave. emergency!!! This seems shaddy, and im sorry. thanks for all of your help though :)

OpenStudy (danjs):

the question is shady? lol.

OpenStudy (anonymous):

Shaddy...

OpenStudy (danjs):

FOr the future, Use those 2 x values in the equation for the line, and get the 2 y values, those are your two intersection points.

OpenStudy (anonymous):

:)